# TS-PEFT: Unveiling Token-Level Redundancy in Parameter-Efficient Fine-Tuning

### v2

@@ Abstract
Current Parameter-Efficient Fine-Tuning (PEFT) methods typically operate under an implicit assumption: once a target module is selected, every token passing through it contributes equally to the downstream task and requires a parameter update. In this paper, we challenge this convention and unveil a pervasive token-level redundancy in the fine-tuning of large models. We propose TS-PEFT, a theoretically grounded framework utilizing proximal optimization to dynamically identify and skip redundant token updates during training. Our extensive experiments across Natural Language Understanding, Commonsense Reasoning, and Visual Instruction Tuning demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise. Strikingly, by discarding 40%-60% of token updates, TS-PEFT consistently matches or surpasses the performance of dense baselines (e.g., LoRA, DoRA). Furthermore, we provide an in-depth analysis revealing that the learned token-level sparsity serves as a superior indicator of module importance compared to traditional weight norms, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
@@ 1 Introduction
Large models (LMs) have demonstrated remarkable capabilities across Natural Language Processing (NLP) yao2024survey and Computer Vision (CV) croitoru2023diffusion; liu2024sora. To adapt these massive models to downstream tasks efficiently, Parameter-Efficient Fine-Tuning (PEFT) has become the de facto paradigmhu2023llm; hanparameter. Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model. These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
However, despite their distinct architectural designs, these methods share a common, implicit assumption: the ”Dense Update” assumption. That is, once a module (e.g., a self-attention layer) is targeted for fine-tuning, the learnable modification is indiscriminately applied to every token position in the input sequence. Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
@@
In this paper, we challenge this convention and investigate a fundamental scientific question: Is it necessary to apply fine-tuning updates to all token positions? Given the strong representational power of pre-trained backbones, we hypothesize that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pre-trained weights. Under this hypothesis, the standard dense update strategy is not only computationally superfluous but may also introduce optimization noise and interference, potentially hindering model performance.
@@
Recent works like CFT ruan2025enhancing have demonstrated that masking out non-critical tokens during training improves reasoning. However, CFT treats redundancy as a static data filtering problem, requiring costly offline counterfactual perturbations to generate fixed masks. In contrast, to systematically verify this hypothesis and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT). Unlike prior approaches, TS-PEFT serves as a learnable probe that dynamically decides whether to update a token or keep its representation frozen. We formulate this selection process as a sparsity-regularized optimization problem, employing proximal optimization and an Adam-style adaptive thresholding mechanism to ensure stable, end-to-end training of the gating decisions.
Our empirical findings unveil a pervasive token-level redundancy in standard PEFT. Across diverse benchmarks including Commonsense Reasoning, Visual Instruction Tuning, and Natural Language Understanding, we observe that discarding 40%-60% of token updates does not degrade performance. Strikingly, TS-PEFT often outperforms its dense counterparts (e.g., LoRA, DoRA) by explicitly filtering out redundant updates. While current hardware constraints on sparse operations limit immediate wall-clock speedups, our work establishes a theoretical computational bound, demonstrating that nearly half of the fine-tuning computation in current paradigms is effectively unnecessary.
@@
Furthermore, our analysis exposes a profound connection between token sparsity and module importance. We discover that the learned sparsity pattern is not random; modules that exhibit low sparsity (i.e., require frequent updates) are significantly more critical to the downstream task than those with high sparsity. This insight allows token-level sparsity to serve as a superior indicator for module selection compared to traditional weight-norm-based metrics.
In summary, our contributions are as follows:
@@
  • Unveiling Redundancy: We identify and quantify that nearly 40%-60% of token updates in standard PEFT are redundant. We empirically demonstrate that indiscriminately updating these tokens is suboptimal and introduces optimization noise.
@@
  • Methodological Framework: We propose TS-PEFT, a theoretically grounded framework that employs proximal optimization to dynamically learn token-selective policies in an end-to-end manner, bypassing the need for costly offline heuristics.
@@
  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
@@ 2 Related Work
Low-Rank Adaptation. Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while preserving downstream performance. LoRA hu2022lora introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models; after training, these matrices can be merged back into the base model without incurring additional inference cost. Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead. However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update. That is to say, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension. This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
@@
Sparsity. Sparsity-guided methods have been widely used to improve the efficiency and interpretability of large language models. By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes. Mixture-of-experts (MoE) models, exemplified by Switch Transformers fedus2022switch and Mixtral 8x7B jiang2024mixtral, employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters, and several recent works have combined MoE architectures with LoRA-style adapters to further enhance performance chen2024llava; liu2024moe; zadouripushing . Another typical use of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets ma2023llm; xubesa; xiasheared. tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
@@
We observe that these methods exploit redundancy at different levels. MoE-based approaches typically target modular-level sparsity, whereas pruning focuses on parameter-level sparsity. In contrast, our work explores token-level sparsity within the context of PEFT, introducing a gating mechanism to selectively skip updates for tokens requiring no task-specific adaptation. Therefore, our method is independent of these architectures and can be theoretically plugged in as a detachable module on top of LoRA-MoE, further eliminating redundant computations.
@@ 3 Method
This section presents the overall theoretical framework and optimization process of TS-PEFT. We first formalize the LoRA-style PEFT procedure and introduce a token-level control function $\operatorname{S}$ to determine whether a PEFT update should be applied at each position. Then, we formulate the learning of the scalar threshold $\tau$ in $\operatorname{S}$ as an optimization problem with a sparsity regularization term. Next, to address the non-differentiability of $\operatorname{S}$ with respect to $\tau$ and the issue that gradients are nearly zero at most positions, we derive a computable approximate gradient. Finally, we combine this approximate gradient with momentum and adaptive learning-rate mechanisms so that it can be seamlessly integrated into the Adam optimizer, and we provide the overall algorithmic description of the TS-PEFT training process.
@@ 3.1 Token-level Selective PEFT Formulation
Here, $\operatorname{S}(\cdot)\in\{0,1\}$ serves as a gating variable: when $\operatorname{S}=1$, the update is enabled at that position, and when $\operatorname{S}=0$, the position retains the base model output. To instantiate Eq.(2), we focus on the relative magnitude $r_{i}$ of the PEFT update with respect to the base model output at each position $i$: $r_{i}=\|\operatorname{M}(x_{i})\|_{2}/\|W_{0}x_{i}\|_{2}$. This is because the feature norm values at different layers vary significantly. The relative values enable each layer to adaptively adjust according to its own weight change ratio. Intuitively, a larger $r_{i}$ means that the PEFT update induces a stronger perturbation relative to the base model output at that position and is more likely to be task-relevant, whereas a small $r_{i}$ indicates that the modification is minor and can be treated as negligible. Therefore, we introduce a learnable scalar threshold $\tau$ for each target layer and define the control function accordingly:
@@
In other words, when the relative magnitude at a given position is not lower than $\tau$, we consider the update at that position worth retaining; otherwise, the update for that position is skipped. This results in a token-level sparse update pattern along the sequence dimension within each selected PEFT module, with the threshold $\tau$ directly controlling the overall sparsity level.
However, this threshold-based step gating introduces two training challenges: first, $\operatorname{S}$ is a non-smooth step function with respect to $\tau$, so standard backpropagation provides almost no useful gradients; second, we want $\operatorname{S}$ to activate PEFT updates only on a subset of token positions while keeping the remaining ones identical to the backbone, which requires explicit control of the overall sparsity level during threshold learning. The next subsections tackle these two issues via a proximal objective and a computable gradient approximation.
@@ 3.2 Proximal Optimization for the Threshold
Due to the training difficulties mentioned above, directly using first-order optimization methods cannot effectively update $\tau$. Therefore, to ensure that the threshold can be stably updated during training, we formulate its learning process as a proximal optimization problem, allowing it to follow the changes in the original task loss while explicitly favoring a sparser update pattern. Specifically, our objective function is defined as follows:
@@
where $\ell(\tau)$ denotes the loss of the target task, and the second term penalizes the number of excessively activated tokens to encourage a sparser update, $\mathbbm{1}(\cdot)$ is the indicator function, and $\lambda$ serves as a hyperparameter that balances the sparsity level.
However, since the indicator function is non-differentiable, we adopt proximal optimization techniques and introduce a quadratic regularization term in the threshold update to limit its fluctuation parikh2014proximal. Given the threshold $\tau^{k}$ at the $k$-th iteration, the update $\tau^{k+1}$ can be written as:
$\underset{\tau}{\mathrm{argmin}}\left[\ell(\tau)+\frac{\partial\ell}{\partial\tau}(\tau-\tau^{k})+\frac{1}{2\alpha_{k}}\left|\tau-\tau^{k}\right|^{2}+\lambda P(\tau)\right],$ (5)
@@
To solve Eq.(5), we need to compute $\partial\ell/\partial\tau$ and $\partial P/\partial\tau$. Thus, we first expand the output in Eq.(2) along the hidden dimension. Let the hidden representation of the $i$-th token at dimension $j\in\{0,\dots,H-1\}$ be denoted as:
@@ 3.3 Gradient Approximation for the Step Function
In the previous subsection, we obtained the derivative expressions of the loss function and the sparsity penalty term with respect to the threshold $\tau$. However, all these derivatives depend on the partial derivative of the step gating $\mathbbm{1}(r_{i}\geq\tau)$ with respect to $\tau$, whose numerical value is almost zero everywhere, making it impossible to directly perform gradient updates. Therefore, to enable stable learning of the threshold through backpropagation, this section constructs a computable gradient approximation (based on the perspective of proximal optimization) to replace the non-differentiable derivative of the indicator function.
@@
A straightforward approximation is to treat the gradient of the step function as a constant s,s>0-s,s>0, that is:
@@
Substituting this approximation into Eq.(7)-(9) yields the “rough” gradients of the loss term and the sparsity term with respect to the threshold:
@@
However, we observe that Eq.(12) will always push $\tau$ upward, while the direction of Eq.(11) depends on the sign of $\mu_{i}$, making Eq.(10) a rather coarse approximation. To avoid unstable oscillation, we further introduce a consistency mask so that only tokens whose PEFT updates “should be activated” contribute to the threshold update. Concretely, we only use positions where the current gating decision agrees with their importance estimate ($\mathbbm{1}(\mu_{i}\geq 0)=\mathbbm{1}(r_{i}\geq\tau)$). Therefore, we revise Eq.(11) and Eq.(12) as follows:
@@
The constraints can be interpreted as follows: whenever $\tau$ is either large enough ($\mu_{i}{\geq}0$, $r_{i}{<}\tau$) or small enough ($\mu_{i}{\leq}0$, $r_{i}{\geq}\tau$) to render the update of $\tau$ unnecessary, Eq.(13) zeros $\partial\ell/\ \partial\tau$; similarly, when $\tau$ is large enough ($r_{i}{<}\tau$, $\lambda$ is a positive constant), Eq.(14) nullifies $\partial P/\ \partial\tau$.

### v3

@@ Abstract
Current Parameter-Efficient Fine-Tuning (PEFT) methods typically operate under an implicit assumption: Once a target module is selected, every token passing through it contributes equally to the downstream task and requires a parameter update. In this paper, we challenge this convention by revealing a pervasive token-level redundancy in the fine-tuning of large models (LMs). We propose TS-PEFT, a theoretical framework utilizing proximal optimization that acts as a dynamic probe to identify token-level redundancy during the fine-tuning process. Extensive experiments demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise. Surprisingly, by discarding 30%-70% of token updates, TS-PEFT consistently matches or exceeds the performance of dense baselines such as LoRA, DoRA. Our in-depth analysis shows that the learned token-level sparsity is a superior indicator of module importance compared to traditional weight criteria, providing a novel data-driven perspective on the intrinsic adaptation mechanism of LMs.
@@ 1 Introduction
Large models (LMs) are demonstrating astonishing capabilities in Natural Language Processing (NLP) Yao et al. (2024); Guo et al. (2025) and Computer Vision (CV) Croitoru et al. (2023); Liu et al. (2024d). To effectively use these large-scale models in downstream tasks, the dominant paradigm is Parameter-Efficient Fine-Tuning (PEFT)Hu et al. (2023); Han et al. (2024). In these approaches, prompt-based methods such as Prefix-Tuning Li and Liang (2021) and Prompt-Tuning Lester et al. (2021) train additional input tokens to change model behavior and low-rank adaptation (LoRA) Hu et al. (2022) and its variants Zhang et al. (2023); Liu et al. (2024c, c) change internal representations by injecting a small number of trainable parameters or corridors, i.e. low-rank components, into the base model. These methods effectively decrease trainable parameters while maintaining similar, or comparable, performance to full parameter fine-tuning.
However, despite their distinct architectural designs, these methods share a common and implicit assumption that they all employ dense updates. That is, once a module like a self-attention layer is used for fine-tuning, the learnable modification is applied indiscriminately to every token position in the input sequence. Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is usually written as:
@@
In this paper, we challenge this convention and investigate a fundamental scientific question: Is it necessary to apply fine-tuning updates to all token positions? Given the strong representational power of pretrained base models, we assume that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pretrained weights. Under this assumption, the standard dense update strategy is suboptimal, primarily because it introduces optimization noise that forces the model to fit gradients on irrelevant tokens, distorting the well-learned features.
@@
Recent works like CFT Ruan et al. (2025) have demonstrated that masking non-critical tokens during training can improve reasoning. This finding corroborates our intuition that, just as imposing a uniform penalty on all tokens in pretraining may hurt generalization, updating all tokens indiscriminately in PEFT similarly introduces redundancy and noise. To systematically verify this assumption and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT). Unlike prior approaches, TS-PEFT serves as a learnable probe that dynamically decides whether to update a token or keep its representation frozen. We formulate this selection process as a sparsity-regularized optimization problem, employing proximal optimization and an Adam-style adaptive thresholding mechanism to ensure stable, end-to-end training of the gating decisions.
Our empirical findings unveil a pervasive token-level redundancy. Across many benchmarks, we observe that nearly 30%-70% of token updates are actually redundant. Crucially, we prove that indiscriminately updating these tokens is detrimental. As evidenced by our analysis, forcing updates on these redundant positions leads to a “performance cliff”, confirming that redundancy in PEFT acts as harmful noise. By filtering out this noise, TS-PEFT consistently matches or exceeds the performance of dense baselines (e.g., LoRA, DoRA) while utilizing significantly fewer active parameters.
@@
Furthermore, our analysis exposes a profound connection between token sparsity and module importance. We discover that the learned sparsity pattern is not random. The modules that exhibit low sparsity, i.e., requiring frequent updates, are significantly more important to the downstream task than those with high sparsity. This insight allows token-level sparsity to serve as a superior indicator for module selection compared to traditional weight-norm-based metrics. In summary, our contributions are as follows:
@@
  • Unveiling Redundancy: We identify and quantify that in standard PEFT, nearly 30%-70% of token updates are redundant. We empirically demonstrate that indiscriminately updating these tokens is suboptimal and introduces optimization noise.
@@
  • Methodological Framework: We propose TS-PEFT, a theoretical framework that employs proximal optimization to dynamically learn token-selective strategies in an end-to-end manner.
@@
  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of LMs.
@@ 2 Related Work
Low-Rank Adaptation. Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while maintaining downstream performance. LoRA Hu et al. (2022) introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models. After training, these matrices can be merged back into the base model without incurring additional inference cost. Based on this concept, AdaLoRA Zhang et al. (2023) adaptively redistributes the effective rank across different layers to better meet task-specific requirements. DoRA Liu et al. (2024c) refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA Kopiczko et al. (2024) utilizes shared low-rank bases to further reduce the parameter overhead for each task. However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update. That is to say, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension. This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
@@
Sparsity. The methods of sparsity guidance have been widely used to improve the efficiency and interpretability of large language models (LLMs). By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes. Mixture-of-experts (MoE) models, such as Switch Transformers Fedus et al. (2022) and Mixtral 8x7B Jiang et al. (2024), employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters. Recently, several works have combined MoE architectures with LoRA-style adapters to further enhance performance Chen et al. (2024); Liu et al. (2024b); Zadouri et al. (2024). Another typical application of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets Ma et al. (2023); Xu et al. (2024); Xia et al. (2024). Tan et al. (2024) further introduces sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
@@
We observe that these methods all utilize redundancy at different levels. MoE-based approaches typically target modular-level sparsity, whereas pruning focuses on parameter-level sparsity. In contrast, our work explores token-level sparsity within the context of PEFT, introducing a gating mechanism to selectively skip updates for tokens requiring no task-specific adaptation. Therefore, our approach is not constrained by these architectures, and theoretically it can be directly embedded as a detachable module into all LoRA variants, including LoRA-MoE, thereby further eliminating redundant computations.
@@ 3 Method
This section introduces the overall theoretical framework and optimization process of TS-PEFT. We first formalize the LoRA-style PEFT procedure and introduce a token-level control function $\operatorname{S}$ to determine whether a PEFT update should be applied at each position. Then, we formulate the learning of the scalar threshold $\tau$ in $\operatorname{S}$ as an optimization problem with a sparsity regularization term. Next, to address the non-differentiability of $\operatorname{S}$ with respect to $\tau$ and the issue that gradients are nearly zero at most positions, we derive a computable approximate gradient. Finally, we combine this approximate gradient with momentum and adaptive learning rate mechanisms so that it can be seamlessly integrated into the Adam optimizer, and we provide the overall algorithmic description of the TS-PEFT training process.
@@ 3.1 Token-level Selective PEFT Formulation
Here, $\operatorname{S}(\cdot)\in\{0,1\}$ serves as a gating variable. When $\operatorname{S}=1$, the update is enabled at that position, while when $\operatorname{S}=0$, the position retains the base model output. To instantiate Eq.(2), we focus on the relative magnitude $r_{i}$ of the PEFT update with respect to the base model output at each position $i$: $r_{i}=\|\operatorname{M}(x_{i})\|_{2}/\|W_{0}x_{i}\|_{2}$. This is because the feature norm values at different layers vary significantly. The relative values enable each layer to adaptively adjust according to its own weight change ratio. Intuitively, a larger $r_{i}$ means that the PEFT update induces a stronger perturbation relative to the base model output at that position and is more likely to be relevant to the task, whereas a small $r_{i}$ indicates that the modification is minor and can be treated as negligible. Therefore, we introduce a learnable scalar threshold $\tau$ for each target layer and define the control function accordingly:
@@
In other words, if the relative magnitude at a certain position is not lower than $\tau$, we consider the update at that position worth retaining. Otherwise, the update for that position is skipped. This will form a token-level sparse update pattern along the sequence dimension within each selected PEFT module, and the threshold $\tau$ controls the overall sparsity level.
However, this threshold-based step gating introduces two training challenges: First, $\operatorname{S}$ is a non-smooth step function with respect to $\tau$, so standard backpropagation provides almost no useful gradients. Second, we want $\operatorname{S}$ to activate PEFT updates only on a subset of token positions while keeping the remaining ones identical to the base model. This requires explicit control of the overall sparsity level during threshold learning. The next subsections tackle these two issues via a proximal objective and a computable gradient approximation.
@@ 3.2 Proximal Optimization for the Threshold
Due to the aforementioned training difficulties, directly using first-order optimization methods cannot effectively update $\tau$. Thus, in order to ensure that the threshold can be updated stably during the training process, we formulate its learning process as a proximal optimization problem, enabling it to follow the changes in the original task loss and explicitly favor a sparser update pattern. Specifically, our objective function is defined as follows:
@@
where $\ell(\tau)$ denotes the loss of the target task, and the second term penalizes the number of excessively activated tokens to encourage a sparser update. $\mathbbm{1}(\cdot)$ is an indicator function, and $\lambda$, as a hyperparameter, is used to balance the sparsity level.
We employ proximal optimization to derive the update rule for parameter $\tau$ within a stochastic gradient descent (SGD) framework Parikh et al. (2014). Given the threshold $\tau^{k}$ at the $k$-th iteration, the update $\tau^{k+1}$ can be written as:
$\underset{\tau}{\mathrm{argmin}}\left[\ell(\tau^{k})+\frac{\partial\ell}{\partial\tau}(\tau-\tau^{k})+\frac{1}{2\alpha_{k}}\left|\tau-\tau^{k}\right|^{2}+\lambda P(\tau)\right],$ (5)
@@
To solve Eq.(5), we need to compute $\partial\ell/\partial\tau$ and $\partial P/\partial\tau$. Thus, we first expand the output of Eq.(2) along the hidden dimension. Let the hidden representation of the $i$-th token at dimension $j\in\{0,\dots,H-1\}$ be denoted as:
@@ 3.3 Gradient Approximation for the Step Function
In the previous subsection, we obtain the derivative expressions of the loss function and the sparsity penalty term with respect to the threshold $\tau$. However, all these derivatives depend on the partial derivative of the step gating $\mathbbm{1}(r_{i}\geq\tau)$ with respect to $\tau$, and its numerical value is almost zero everywhere, making it impossible to directly perform gradient updates. Therefore, to enable stable learning of the threshold through backpropagation, this section constructs a computable gradient approximation based on the perspective of proximal optimization to replace the non-differentiable derivative of the indicator function.
@@
A straightforward approximation, inspired by Bengio et al. (2013), is to treat the gradient of the step function as a constant s,s>0-s,s>0, that is:
@@
Substituting this approximation value into Eq.(7)-(9) yields the rough gradients of the loss term and the sparsity term with respect to the threshold:
@@
Eq.(10) is a rather rough approximation, which causes highly oscillatory updates of $\tau$ and results in poor model performance. We believe this is because Eq.(10) should be 0 for certain value domains, but for the update, we set it to s-s. Therefore, to limit this effect, we add indicator conditions. We modify Eq.(11) and Eq.(12) as follows:
@@
The constraints can be understood as follows: Whenever $\tau$ is either large enough ($\mu_{i}{\geq}0$, $r_{i}{<}\tau$) or small enough ($\mu_{i}{\leq}0$, $r_{i}{\geq}\tau$) to render the update of $\tau$ unnecessary, Eq.(13) zeros $\partial\ell/\ \partial\tau$. Similarly, when $\tau$ is large enough ($r_{i}{<}\tau$, $\lambda$ is a positive constant), Eq.(14) nullifies $\partial P/\ \partial\tau$.